Prove that the dot product between two vectors is commutative and not associative


1
Expert's answer
2021-06-15T10:37:17-0400

Assume two vectors "\vec u=x \hat i+y\hat j+z\hat k, \vec v=a \hat i+b\hat j+c\hat k" .


Now, "\vec u.\vec v=(x \hat i+y\hat j+z\hat k).(a \hat i+b\hat j+c\hat k)""=xa+yb+zc" ...(i)


And "\vec v.\vec w=(a \hat i+b\hat j+c\hat k).(x \hat i+y\hat j+z\hat k)""=ax+by+cz" ...(ii)


From (i) and (ii), "\vec u.\vec v=\vec v.\vec u"


Hence, commutative.


Also, assume another vector "\vec w=p \hat i+q\hat j+r\hat k"

"(\vec u.\vec v).\vec w=[(x \hat i+y\hat j+z\hat k).(a \hat i+b\hat j+c\hat k)].(p \hat i+q\hat j+r\hat k)
\\=[xa+yb+zc].(p \hat i+q\hat j+r\hat k)"

which is not defined as dot product of a scalar and vector quantity is not defined.


Now,

"\vec u.(\vec v.\vec w)=(x \hat i+y\hat j+z\hat k).[(a \hat i+b\hat j+c\hat k).(p \hat i+q\hat j+r\hat k)]
\\=(x \hat i+y\hat j+z\hat k).(ap+bq+cr)"

which is not defined as dot product of a scalar and vector quantity is not defined.


Hence, not associative.